OpenStudy (anonymous):

2x^2 + 39 = 57 solve the equation using square roots.

OpenStudy (anonymous):

Change the equation so you get \[x^2 = number\] and then take the square root of both sides

OpenStudy (anonymous):

@tomo

OpenStudy (anonymous):

2x^2 + 39 = 57 2x^2 = 57-39 2x^2 = 18 x^2 = 18/2 x^2 = 9 x = sqrt(9) // remember there is a positive and negative value for the square root. post what you get for x.